OpenStudy (anonymous):

use the discriminant to determine the type and number of solutions. -3x^2+2x+4=0 A) 2 rational solutions B) 2 imaginary solutions C) 1 rational solution D) 2 irrational solutions

OpenStudy (jhannybean):

The discriminant is \[\large b^2-4ac\]

OpenStudy (jhannybean):

\[\large b = 2 \ , \ a = -3 \ , \ c = 4 \]

OpenStudy (luigi0210):

If it's positive/perfect square it has 2 real solutions If it's 0 it has one repeating solution If it's negative it has no real solutions/2imaginary solutions

OpenStudy (jhannybean):

\[\large b^2-4ac = (2)^2-4(-3)(4) = ?\]

OpenStudy (luigi0210):

Smartypants

OpenStudy (jhannybean):

Ty :)

OpenStudy (anonymous):

i get b^2-4ac=52

OpenStudy (luigi0210):

Good

OpenStudy (anonymous):

so its C?

OpenStudy (jhannybean):

Nope. it's a positive outcome, hence what @Luigi0210 stated.